# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ Abstract
On the even-circumference geometries $L_{y}=4$ and $6$, we reproduce the weak crossover, with a correlation length of $\sim 2$ lattice sites, found by Kuhlenkamp et al. for $L_{y}=6$ and $\Phi_{\text{ext}}=0$ [11]. This is consistent with the absence of translation symmetry-breaking in the even-circumference Mott limit. Upon threading $\Phi_{\text{ext}}=\pi/2$ external flux through each of these systems, we observe dramatically larger values of the correlation length for $U$ near $U_{c}$ (Fig. 3), notably in the same neutral sector in which the odd-circumference correlation length diverges [see Fig. 3(c,d)]. We also perform an extrapolation to infinite bond dimension following Ref. [52]; see SM for details [29]. The maximal extrapolated value is $\xi=6.6$ at $L_{y}=4$, increasing substantially to $\xi=9.1$ at $L_{y}=6$. Consistent with the view that accurate estimation of the 2+1D gap and other properties requires consideration of all threaded cylinder fluxes [53], we conjecture that the $\Phi_{\text{ext}}=\pi/2$ threaded system is more reflective of the thermodynamic limit than the unthreaded system, and that larger even-$L_{y}$ geometries will exhibit an increasingly-sharp sequence of crossovers toward 2+1D critical behavior in the charge-neutral sector.
Figure 4: Density-density [left panel] and spin-spin [right panel] connected correlations for odd-circumference cylinders with $U$ tuned to the respective critical points, $U_{*}(L_{y}=3)=11.1$ and $U_{*}(L_{y}=5)=11.6$. Both $L_{y}=5$ correlation functions are scaled down by 100 for ease of viewing. (a) Density-density correlations decay nearly algebraically at the IQH-CSL transition, pointing to the presence of gapless charge-neutral excitations in the 2+1D limit. (b) In contrast, correlations of the spin density decay exponentially with no apparent enhancement from $L_{y}=3$ to $5$, suggesting that the excitations are spin-rotationally-invariant singlets.
Remarkably, the critical fluctuations at the odd-$L_{y}$ critical points manifest not only in the order parameter, but also in the charge density. This is seen manifestly in the connected density-density correlations $\langle n(x)n(0)\rangle$, with $n(x)$ averaged over the ring $x$, which we find to decay approximately algebraically, i.e., linearly on a log-log scale [see Fig. 4(a)] However, estimation of the 2+1D critical exponents from these data is hampered by the fact that the critical features of the thermodynamic limit need only show up in the regime $1\ll x\ll L_{y}$, which is inaccessible for the presently-tractable system sizes. While PH-glide is no longer broken for $L_{y}=4,6$ and the operator Eq. (3) no longer condenses (see Fig. 3a), we also observe that their enhanced correlation length is accompanied by an apparent algebraic plateau in $\langle n(x)n(0)\rangle$ over the scale of the correlation length (see SM [29]).
Curiously, we do not observe critical behavior in the odd-$L_{y}$ connected correlations of the spin density $S_{z}$ [Fig. 4(b)], which is also neutral under the conserved charge and spin $U(1)$ symmetries. However, since $S_{z}$ may be rotated into $S_{x}$ by a global spin rotation, under which the CSL is invariant, this is consistent with our not observing a diverging correlation length in the $(Q,S_{z})=(0,\hbar)$ sector of the transfer matrix. Altogether, under the assumption that this behavior persists for the entire sequence of odd circumferences, our observations suggest that the 2+1D critical point hosts gapless, charge-neutral spin-singlet excitations.
We remark that these data are consistent with the partonic view that the transition into the IQH corresponds to chargon condensation. While the chargon is not gauge invariant and may not develop conventional long-range connected correlations, at the critical point the combination of the chargon, anti-chargon and gauge string is a gapless neutral excitation [42]. Interestingly, the power law correlations presented in Fig. 4 show that the physical charge density operator, which has nonzero overlap with the gauge-invariant chargon density $b^{\dagger}(x)b(x)$, produces a state with a nonzero overlap with the critical excitations when applied to the ground state, raising the intriguing possibility that the critical fluctuations can be experimentally observed.
Discussion—By combining numerical data on both even and odd-circumference cylinder geometries, we have provided evidence that the CSL-IQH transition is in fact continuous in the Hofstadter-Hubbard model at $\pi/2$ flux per triangle and that critical properties are already manifest at the numerically-accessible circumferences $3\leq L_{y}\leq 6$. Our finding that critical fluctuations appear in the charge density correlation function, which may be probed by experimental or numerical measurement of the structure factor, raise the possibility of experimental and numerical investigation of this transition.
This paradigmatic Hofstadter-Hubbard system provides a uniquely minimal and experimentally-relevant setting [10, 11] in which to study a fundamental phase transition at the interface of spin liquid and quantum Hall physics. As such, a more complete understanding of this transition is an important task for future work. Quantitatively, our investigation clarifies the possible observable, numerical signatures that may be used to probe the phase transition in a variety of numerical methods, such as AFQMC, VMC, or fermionic PEPS, which may be able to access larger system sizes or spectral signatures of spin-charge separation [54, 55, 56].
@@
An important broader message of this work is that the ability to extrapolate 2+1D physics from dimensionally-reduced numerical data—namely on infinite cylinder geometries—is greatly enhanced by carefully considering the manner in which candidate ground states and their phase transitions respond to dimensional reduction and to external fields that become boundary effects in the thermodynamic limit. Applying such techniques more systematically, it would be valuable to characterize the contexts in which UV features of Mott phases, namely spin liquids, may be leveraged to better characterize their charge fluctuation-induced phase transitions.
@@
Acknowledgements—We acknowledge useful discussions with E. Altman, S. Chatterjee, O. Gauthé, M. Knap, C. Kuhlenkamp, J.Y. Lee, C. Liu, F. Pollmann, Z. Weinstein, and especially X.Y. Song. The Flatiron Institute is a division of the Simons Foundation. AJM was supported in part by Programmable Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award DE-SC0019443. S.D. acknowledges support from the NSERC PGSD fellowship. M.Z. and S.D. were supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (QSA). This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8683 to T.S. Calculations were performed using the TeNPy Library [57].
@@ Appendix A Hamiltonian construction at all rational fluxes
For the underlying triangular lattice — with the “YC” orientation on the cylinder (where the $\hat{y}$ site-to-site distance is the NN distance $a$) — we take the following Bravais vectors:
[…]
whose corresponding reciprocal lattice vectors are
[…]
SI Choosing the magnetic unit cell
The following discussion applies to both the YC and XC lattice geometries. Let the flux per parallelogram be
@@
Regardless of the parity of $q$, there will be $q$ magnetic sublattice sites in the magnetic unit cell, which we label by $\tau=0,\dots,q-1.$ When $q$ is even, the magnetic Bravais vectors are
@@
whose total area $A_{\text{muc}}$ (for either YC or XC) is precisely that of $2q$ triangles of area $\frac{\sqrt{3}}{4}a^{2}.$ The associated reciprocal lattice vectors are
@@
$\bm{R}_{1}=q\bm{a}_{1}-\frac{q-1}{2}\bm{a}_{2},\qquad\bm{R}_{2}=\bm{a}_{2}\propto\hat{y},,\qquad(q\ \text{odd}),$ (8)
@@
which also satisfy the inverse area relationship, Eq. (7).
In Fig. LABEL:fig:YC_c2zsymm_locations, we illustrate the choice of magnetic unit cell for both odd and even $q$ for the YC triangular lattice boundary conditions. For both even and odd denominator $q,$ note that the enclosing magnetic unit cell contains exactly $2q$ triangles, so $2\pi p\in 2\pi\mathbb{Z}$ total flux. In addition, we position the sublattice sites $\tau$ within the magnetic unit cell in such a way that all the sites within magnetic unit cell $\bm{R}$ transform under $C_{2z}$ into sites in $-\bm{R}$. In particular, the sites in the home magnetic unit cell transform into one another in a clean way. We achieved this by further partitioning the odd-$q$ geometries into two subclasses. In Fig. LABEL:fig:XC_c2zsymm_locations, we make the same specifications for the XC boundary conditions.
SII Hamiltonian and choice of vector potential
Placing the underlying triangular lattice on the cylinder, let the resulting system have circumference $L_{y}$, i.e., the points $\bm{r}\sim\bm{r}+L_{y}\bm{a}_{2}$ are geometrically identified. Let the lattice sites be indexed by
$\bm{r}=x\bm{a}_{1}+y\bm{a}_{2},$ (11)
which defines our coordinate system, where $(x,y)\in\mathbb{Z}^{2}.$ Since $\bm{a}_{2}$ is purely vertical in both geometries, then $x$ indexes the rings, and $y$ is a integer spatial index within each ring. To generate complex hoppings and modify the translation algebra to a magnetic translation algebra, we introduce an external $U(1)$ vector potential. We assume that it is linear and that its curl is the magnetic field:
@@
[…]
@@
To constrain the form Eq. (12), we remark that the terms $x\hat{x}$ and $y\hat{y}$ are curl-free, whereas $Bx\hat{y}$ and $-By\hat{x}$ — the standard Landau gauges on the rectangular lattice — have the correct curl. We therefore reason that the most general permissible $M$ is given by
@@
where $\alpha,\beta,\gamma\in\mathbb{R}$. We will now demand that the choice of vector potential is compatible with our chosen geometry of magnetic Bravais vectors. Specifically, we ask that the Peierl’s hoppings generated by $\bm{A}$ are periodic in the chosen Bravais vectors. This allows us to place the system easily on a torus or cylinder, provided the system tiles an integer number of magnetic unit cells. Quantitatively, we demand that the RHS of Eq. (13) is invariant under shifts of the argument of $\bm{A}$ by the magnetic Bravais vectors $\bm{R}_{1}$ and $\bm{R}_{2}$, for all bond vectors $\bm{r}^{\prime}-\bm{r}$. By linearity of the vector potential, and since the bond vectors are linear combinations of $\bm{a}_{1}$ and $\bm{a}_{2}$, it suffices to require that
@@
for all $i,j$. Following Ref. [58], an elegant solution can be found in terms of the reciprocal vectors of the underlying triangular lattice. We take the following since $\bm{b}_{1}\propto\hat{x}$ for both YC and XC geometries, and we want the vector potential to be independent of $y$:
$\bm{A}(\bm{r})=\frac{\Phi}{(2\pi)^{2}}(\bm{r}\cdot\bm{b}_{1})\bm{b}_{2}.$ (18)
@@
[×$\left[\nabla\times\left((\bm{r}\cdot\bm{Q})\bm{P}\right)\right]_{i}$ $=\epsilon_{ijk}\partial_{j}\left(\bm{r}_{\ell}\bm{Q}_{\ell}\bm{P}_{k}\right)=\epsilon_{ijk}\bm{Q}_{j}\bm{P}_{k}=(\bm{Q}\times\bm{P})_{i},$ (19)
@@
$\bm{B}(\bm{r})$ =×$=\nabla\times\bm{A}(\bm{r})=\frac{\Phi}{(2\pi)^{2}}\bm{b}_{1}\times\bm{b}_{2}=\frac{\Phi}{(2\pi)^{2}}\frac{(2\pi)^{2}}{A_{\text{par}}}\hat{z}=\frac{\Phi}{A_{\text{par}}}\hat{z}.$ (20)
Moreover, it can be shown to have the required periodicity property:

### v2

@@ Abstract
On the even circumference geometries $L_{y}=4,6$ we reproduce the weak crossover, with a correlation length of $\lesssim 1.5$ lattice sites, found by Kuhlenkamp et al. for $L_{y}=6$ and $\Phi_{\text{ext}}=0$ [19] (App. F1). Remarkably, upon threading $\Phi_{\text{ext}}=\pi/2$, the correlation length in the range $10<U<12$ increases dramatically, notably in the same $(Q,S_{z})=0$ sector in which our odd-circumference correlation lengths diverge [see Fig. 3(c,d)]. At $L_{y}=4$, extrapolation to infinite bond dimension following Ref. [64] predicts a maximal extrapolated value of $\xi_{L_{y}=4}^{\chi\to\infty}=6.6$ in units of cylinder rings (see App. F2). While at $L_{y}=6$, $\xi=5.8$ at the largest bond dimension $\chi=32768$, the correlation length is still growing rapidly with $\chi$. This hampers the precise extrapolation of $\xi_{L_{y}=6}^{\chi\to\infty}$ but also indicates it substantially exceeds $\xi_{L_{y}=4}^{\chi\to\infty}$. This suggests that the larger even-$L_{y}$ geometries will exhibit a sequence of increasingly-sharp crossovers toward 2+1D criticality.
We now discuss the behavior of correlation functions. The connected current correlator $C_{JJ}(x)=\langle J_{0}^{\perp}J_{x}^{\perp}\rangle_{\text{c}}$ exhibits power-law decay at the approximate transition points $U_{c}(L_{y}=3)=11.1$ and $U_{c}(L_{y}=5)=11.6$ (Fig. 4a), chosen as $U=U_{c}$ where the correlation length is largest at the largest available $\chi$. The critical fluctuations also manifest in the charge density: the connected correlations $C_{nn}(x)=\langle n_{0}n_{x}\rangle_{\text{c}}$ (with $n_{x}$ averaged over ring $x$) are negative and decay algebraically (Fig. 4b), but fall off much more rapidly than the current correlations. While there is no SSB for $L_{y}=4,6$ we still observe an enhanced correlation length (see Fig. 3) and slow decay of the current and density correlations (App. H).
On the other hand, the connected spin correlations $C_{S^{z}S^{z}}(x)=\langle S^{z}_{0}S^{z}_{x}\rangle_{\text{c}}$ (with $S^{z}_{x}$ averaged over ring $x$) exhibit clear exponential decay at the odd-$L_{y}$ critical points, oscillating in sign with a period of two rings, with no apparent enhancement from $L_{y}=3$ to $L_{y}=5$ (Fig. 4c). The electron correlators $C_{c^{\dagger}c}(x)=L_{y}^{-2}\sum_{k_{y}}|\langle c^{\dagger}(0,k_{y})c(x,k_{y})\rangle|$ likewise decay exponentially (Fig. 4d), so that both spin and electron excitations are gapped at the transition. Should these behaviors persist in the thermodynamic limit, they imply that the 2+1D critical point is continuous and hosts gapless, charge-neutral spin-singlet fluctuations, namely of the current.
Figure 4: Connected correlation functions on the odd $L_{y}$ cylinders at their critical point, $U_{c}(L_{y}=3)=11.1$ (yellow) and $U_{c}(L_{y}=5)=11.6$ (blue). The $L_{y}=5$ correlations are reduced by $\times 10$ for ease of viewing. Upper: Current-current correlations and negative of density-density correlations, approaching quasi-long range order as a function of MPS bond dimension $\chi$ (increasing light to dark). Lower: Spin-spin (plotted with alternating sign $(-1)^{x}$) and electron-hole correlations decay exponentially, demonstrating a spin gap.
We comment briefly on the nature of the odd-$L_{y}$ quasi-(1+1)(1+1)D critical points. As the $1e$ and $2e$ correlators remain gapped (App. G1), the data are incompatible with Luttinger and Luther-Emery liquid criticality [65, 66]. The spontaneous breaking of a $\mathbb{Z}_{2}$ symmetry suggests the transition is in the 2D Ising universality class. The “ionic” Hubbard chain [67, 68, 69] shares symmetries and behavior with the studied odd $L_{y}$ systems [70, 71, 72, 73] and has a Mott transition; while the model has subtleties that have complicated the unambiguous determination of exponents, it is likely that the transition is of the Ising type (see App. G2 for a more detailed discussion). Further investigation into the odd $L_{y}$ sequence of transitions, particularly precisely how they evolve to the 2+1D transition and which critical correlations survive, is needed. The prominence of current-current correlations, namely their slow falloff at both short and long distances, suggests they remain prominent at the 2+1D critical point.
Discussion—Our numerical data on both even and odd-circumference cylinder geometries provides confirming evidence of a CSL phase and indicates that the CSL-IQH transition is in fact continuous in the Hofstadter-Hubbard model at $\pi/2$ flux per triangle. The single-electron and spin gaps remain nonzero across the transition, while the current-current correlation function exhibits power-law behavior at the odd-$L_{y}$ critical points and is also prominent on the even-$L_{y}$ geometries. This raises the possibility of experimental detection of the transition and provides valuable guidance for its further numerical examination. Using quantum Monte Carlo, fermionic PEPS, or neural quantum states, one may be able to access wider systems, yielding confirming evidence of quantum criticality, estimates of critical exponents, and possible signatures of conformal invariance.
@@
A broader message of this work is that the ability to extrapolate 2+1D physics from dimensionally-reduced data is greatly enhanced by analyzing how symmetries, ground states and their phase transitions respond to spatial compactification and flux threading. In particular, the properties of transitions into topologically-ordered phases may generally be illuminated by studying cylinder compactifications in which topological ground state degeneracy manifests as SSB [74, 75]. Leveraging such techniques systematically, it would be valuable to characterize the contexts in which UV features of Mott and quantum Hall phases offer improved understanding of their charge fluctuation-driven quantum phase transitions.
Acknowledgements.
@@
Acknowledgements—We acknowledge useful discussions with E. Altman, S. Chatterjee, O. Gauthé, Y-C. He, M. Knap, C. Kuhlenkamp, J.Y. Lee, C. Liu, F. Pollmann, R. Sahay, R. Verresen, Z. Weinstein, Y-H. Zhang, and especially X-Y. Song and A. Vishwanath. S.D. acknowledges support from the NSERC PGSD fellowship. A.J.M. was supported in part by Programmable Quantum Materials, an Energy Frontier Research Center funded by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award DE-SC0019443. M.Z. and S.D. were supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator (QSA). This research is funded in part by the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF8683 to T.S. Calculations were performed using the TeNPy Library [76, 77, 78]. The Flatiron Institute is a division of the Simons Foundation.
@@ Appendix A Hamiltonian construction at all rational fluxes
SI YC Lattice Parameterization
We first describe two coordinate parameterizations of the underlying triangular lattice, beginning in the 2D limit. In the “YC” parameterization of the 2D plane, the $\hat{y}$ site-to-site distance is the nearest-neighbour distance $a$. This corresponds to the Bravais vectors
$\bm{a}_{1}=a\left(\frac{\sqrt{3}}{2}\hat{x}+\frac{1}{2}\hat{y}\right),\qquad\bm{a}_{2}=a\hat{y}\qquad\text{(YC)},$ (1)
and the corresponding reciprocal lattice vectors
$\bm{b}_{1}=\frac{4\pi}{\sqrt{3}a}\hat{x},\qquad\bm{b}_{2}=\frac{2\pi}{a}\left(-\frac{\hat{x}}{\sqrt{3}}+\hat{y}\right)\qquad\text{(YC)}.$ (2)
This leads to a natural cylinder compactification: one may identify each point
$\bm{r}=x\bm{a}_{1}+y\bm{a}_{2}$ (3)
with points $\bm{r}^{\prime}$ for which $y\equiv y^{\prime}\ (\text{mod }L_{y})$ and also $x=x^{\prime}$. We refer to $L_{y}$ as the “circumference” of the cylinder. Doing so for the YC lattice, one obtains the canonical “YC-$L_{y}$” cylinder geometry [8], whose circumference is $L_{y}a$.
SII Choosing the magnetic unit cell
Let the flux per parallelogram (i.e., two triangles) be
@@
Regardless of the parity of $q$, there will be $q$ magnetic sublattice sites in the magnetic unit cell. When $q$ is even, the magnetic Bravais vectors are
@@
whose total area $A_{\text{muc}}$ is precisely that of $2q$ triangles of area $\frac{\sqrt{3}}{4}a^{2}.$ The associated reciprocal lattice vectors are
@@
$\bm{R}_{1}=q\bm{a}_{1}-\frac{q-1}{2}\bm{a}_{2},\qquad\bm{R}_{2}=\bm{a}_{2}\propto\hat{y},\qquad(q\ \text{odd}),$ (9)
@@
which also satisfy the inverse area relationship, Eq. (8). For both even and odd denominator $q,$ note that the enclosing magnetic unit cell contains exactly $2q$ triangles, so $2\pi p\in 2\pi\mathbb{Z}$ total flux.
SIII Hamiltonian and choice of vector potential
In all cases, we adopt the standard convention for the on-site Hubbard repulsion:
$H_{\mathrm{int}}=U\sum_{\bm{r}}n_{\bm{r}\uparrow}n_{\bm{r}\downarrow},\qquad U\geq 0.$ (12)
Now we discuss the hoppings. Placing the underlying triangular lattice on the cylinder as described in Sec. SI, let the resulting system have circumference $L_{y}$, i.e., the points $\bm{r}\sim\bm{r}+L_{y}\bm{a}_{2}$ are geometrically identified. Let the lattice sites be indexed as in Eq. (3), which defines a set of coordinate coefficients $(x,y)\in\mathbb{Z}^{2}.$ Since $\bm{a}_{2}$ is purely vertical, then $x$ indexes the cylinder “rings”, and $y$ is an integer-valued spatial index within each ring. To generate complex hoppings and modify the translation algebra into a magnetic translation algebra, we introduce an external $U(1)$ vector potential. We assume that it is linear and that its curl is the magnetic field:
@@
$t_{\bm{r}^{\prime}\leftarrow\bm{r}}=t\exp\left(i\int^{\bm{r}^{\prime}}_{\bm{r}}\bm{A}(\bm{\ell})\cdot d\bm{\ell}\right)=t\exp\left(i\bm{A}\left(\frac{\bm{r}^{\prime}+\bm{r}}{2}\right)\cdot(\bm{r}^{\prime}-\bm{r})\right),$ (14)
@@
To constrain the form of the vector potential, we remark that the terms $x\hat{x}$ and $y\hat{y}$ are curl-free, whereas $Bx\hat{y}$ and $By\hat{x}$ (the standard Landau gauge potentials on the rectangular lattice) have the correct curl. We therefore reason that the most general permissible $M$ in Eq. (13) is given by
@@
where $\alpha,\beta,\gamma\in\mathbb{R}$. We will now demand that the choice of vector potential is compatible with our chosen geometry of magnetic Bravais vectors. Specifically, we ask that the Peierl’s hoppings generated by $\bm{A}$ are periodic in the chosen magnetic Bravais vectors. This allows us to place the system easily on a torus or cylinder, provided the system is an integer tiling of magnetic unit cells. Quantitatively, we demand that the RHS of Eq. (14) is invariant under shifts of the argument of $\bm{A}$ by the magnetic Bravais vectors $\bm{R}_{1}$ and $\bm{R}_{2}$, for all bond vectors $\bm{r}^{\prime}-\bm{r}$. By linearity of the vector potential, and since the bond vectors are linear combinations of $\bm{a}_{1}$ and $\bm{a}_{2}$, it suffices to require that
@@
for all $i,j$. Following Ref. [79], an elegant solution can be found in terms of the reciprocal vectors of the underlying triangular lattice. We take the following since $\bm{b}_{1}\propto\hat{x}$, and we want the vector potential to be independent of $y$:
$\bm{A}(\bm{r})=\frac{\Phi}{(2\pi)^{2}}(\bm{r}\cdot\bm{b}_{1})\bm{b}_{2},$ (19)
@@
$\left[\nabla\times\left((\bm{r}\cdot\bm{Q})\bm{P}\right)\right]_{i}$ $=\epsilon_{ijk}\partial_{j}\left(\bm{r}_{\ell}\bm{Q}_{\ell}\bm{P}_{k}\right)=\epsilon_{ijk}\bm{Q}_{j}\bm{P}_{k}=(\bm{Q}\times\bm{P})_{i},$ (20)
@@
$\bm{B}(\bm{r})$ $=\nabla\times\bm{A}(\bm{r})=\frac{\Phi}{(2\pi)^{2}}\bm{b}_{1}\times\bm{b}_{2}=\frac{\Phi}{(2\pi)^{2}}\frac{(2\pi)^{2}}{A_{\text{par}}}\hat{z}=\frac{\Phi}{A_{\text{par}}}\hat{z}$. (21)
Moreover, it can be shown to have the required periodicity property. Note that
@@
Specializing to the case $(p,q)=(1,2)$ relevant to the main text, we plot the resulting hopping amplitudes Eq. (16) in Fig. S1(a) for the YC parameterization of the 2D lattice. The hoppings only take values in $\{\pm 1,\pm i\}$ and are evidently periodic in both $\bm{R}_{1}$ and $\bm{R}_{2}$. The color of each bond denotes the complex phase of $t_{\bm{r}^{\prime}\leftarrow\bm{r}}$ (see Eq. (15)) in the direction of the arrow (pictured at the bond centers).
Figure S1: (a) Graphical depiction of the complex phases of the nearest-neighbor hopping amplitudes of Eq. (16) for the $(p,q)=(1,2)$ two-band Hamiltonian in a portion of the infinite 2D plane, where the orientation of the plane is consistent with the YC coordinate system specified by Eqs. (1,6). The color of each bond denotes the complex phase of $t_{\bm{r}^{\prime}\leftarrow\bm{r}}$ (see Eq. (15)) in the direction of the arrow (pictured at the bond centers). The two orange dots indicate the two sites in the magnetic unit cell. Note that the top/bottom and right/left edges are not implicitly identified. (c) Graphical depiction of the fluxes through adjacent loops of a YC-$L_{y}$ cylinder, defined in Sec. SI. The outside surface of each cylindrical segment has a total area of $2L_{y}$ triangles, each having $\pi/2$ flux [pictured]. When $L_{y}$ is odd [upper scenario], this leads to a $\pi$-staggering of the flux through each cylinder ring due to Gauss’ law. When $L_{y}$ is even [lower scenario], the flux through each ring is the same.